how many natural numbers smaller than 10^4 and divisible by 4 can be formed using 01235.


 

Grade:

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Shiwani

One digit number- there will be no one digit number

two digit numbers

12,20,32,52 total 4

3 digit number

we have to fix last two digit for the divisibility by 4

last two digits can be 12,20,32,52 corresponding to these first digit can be chosen in 4 ways leaving 0 out of 5 numbers 0,1,2,3,5

so there are 4*4=16 numbers

 

4 digit numbers

again last two digits will be 12,20,32,52 and first digit can be chosen in 4 ways (leaving 0) and 2nd digit can be chosen in 5 ways so for each there will be 4*5=20 numbers total for all the 12,20,32,52 will be 4*20=80 numbers

 

there will be no 5 digit number

 

so total numbers will be= 4+16+80= 100 numbers

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..